### v1

# Learning Diverse Robot Striking Motions with Diffusion Models and Kinematically Constrained Gradient Guidance

Abstract
Advances in robot learning have enabled robots to generate skills for a variety of tasks. Yet, robot learning is typically sample inefficient, struggles to learn from data sources exhibiting varied behaviors, and does not naturally incorporate constraints. These properties are critical for fast, agile tasks such as playing table tennis. Modern techniques for learning from demonstration improve sample efficiency and scale to diverse data, but are rarely evaluated on agile tasks. In the case of reinforcement learning, achieving good performance requires training on high-fidelity simulators. To overcome these limitations, we develop a novel diffusion modeling approach that is offline, constraint-guided, and expressive of diverse agile behaviors. The key to our approach is a kinematic constraint gradient guidance (KCGG) technique that computes gradients through both the forward kinematics of the robot arm and the diffusion model to direct the sampling process. KCGG minimizes the cost of violating constraints while simultaneously keeping the sampled trajectory in-distribution of the training data. We demonstrate the effectiveness of our approach for time-critical robotic tasks by evaluating KCGG in two challenging domains: simulated air hockey and real table tennis. In simulated air hockey, we achieved a 25.4% increase in block rate, while in table tennis, we saw a 17.3% increase in success rate compared to imitation learning baselines.
I Introduction
Controlling robot motion for agile athletic tasks, such as throwing, catching, or striking a ball, presents substantial challenges. These activities require algorithms for perception, prediction, and motor control, as well as robust hardware designed to handle dynamic actions. Moreover, these tasks demand high levels of precision, where minor errors in decision-making can lead to immediate failures. Recent learning-based approaches for agile tasks are starting to rival hand-tuned control methods. Yet, specifying task execution within these learning-based frameworks often presents challenges. For reinforcement learning (RL) methods, expert designers must define custom reward functions. Although inverse reinforcement learning [1, 2] and adversarial imitation learning methods [3] show potential, they frequently necessitate extensive hyperparameter tuning. In this paper, we approach the problem with imitation learning using a limited and diverse set of demonstrations.
RL and Learning from Demonstration (LfD) are the predominant robot learning methods used to solve agile tasks. Recent works in RL for robotics have tried to improve sample-efficiency [4] with advancements in sim-to-real transfer [5, 6]. However, these works rely on high fidelity simulators [7, 8, 9, 10, 11, 12, 13] and require experts to specify reward or cost functions which may not capture the true expert’s intentions [14]. Similarly, LfD [15, 16, 17, 2, 18] techniques struggle to account for behavioral variance in multi-task settings. Recently, Transformer-based LfD methods such as ACT [19] have drastically improved the sample efficiency of deep learning-based LfD. Both RL and LfD methods do not have the capability to enforce other constraints at test-time dynamically. RL policies assume a fixed MDP formulation and cannot adapt to new reward functions or task constraints. Deep learning-based LfD algorithms also typically do not have a formulation to incorporate novel kinematic or dynamic constraints at test time.
Figure 1: Overview of KCGG: Our diffusion model was trained on four kinesthetically demonstrated strokes. At inference, for each denoising step, our KCGG method computes the cost given the constraint function and intermediate trajectory. Then, the gradient with respect to the prior trajectory through the diffusion model is computed to update the sample.
In this paper, we develop a novel imitation learning based approach that enables robots to learn from small offline datasets to perform agile tasks in dynamical environments. We ground our work in application to striking motions evaluated in established challenge domains for robotics: a virtual air hockey domain [20, 9] and a physical table tennis domain [21, 22]. We show that our diffusion-based approach can learn to condition trajectories to meet environment constraints and perform tasks difficult to specify (e.g. hitting a topspin, a high lob, etc) with just 15 kinesthetic demonstrations per stroke type, which classical motion planning cannot achieve without explicit task heuristics. The key to our approach is the novel formulation of a kinematic constraint gradient guidance (KCGG) technique that enables the robot to balance between constraints and adhering to the demonstration distribution.
Contributions: Our key contributions are:
  • We demonstrate our sampling method can reproduce distinct, multimodal behavior with a diffusion model trained from limited expert demonstrations ($<120$).
  • We propose KCGG, a novel state-of-the-art constraint sampling method for diffusion models. We show that KCGG outperforms baseline diffusion sampling methods in a simulated AirHockey domain. Our method achieves a 21.7% increase in successful blocks.
  • We show KCGG improves a real-world table tennis robot’s ability to return ping-pong balls. Our method improves success rate by 124.5% over diffusion baselines [23, 24] and 17.3% over state of the art imitation learning baselines [19], demonstrating efficacy in agile robotic tasks.
II Related Work
We briefly review prior learning methods for robotic striking motions and key literature on diffusion models.
Learning Striking Motions
LfD [15, 16, 17] aims to enable robots to acquire new skills by learning to imitate an expert. MP-based LfD has been explored extensively for agile striking motions in table tennis by learning movement primitives. For example, Mixture of Movement Primitives (MoMP) [22, 25] learns to select a dynamic movement primitive (DMP) from a library of pre-trained DMPs to generate joint trajectories for returning the ball. Probabilistic movement primitives (ProMP) [26], which model the variability in movements as trajectory distributions, were adapted to generate striking motions that maximize the likelihood of ball return [27, 28, 29, 30]. Movement primitive approaches are sample-efficient; however, generating diverse motions (e.g. slice vs. lob in table tennis) is difficult without a hierarchical approach.
Recently, RL methods that are utilizing heavy computational resources are enabling systems to explore in task space without using demonstrated expert knowledge [21, 31, 32, 33, 34]. However, most of them require large training datasets [21, 32], carefully shaped reward [21], or human behavior modeling [34] which require a simulation. In contrast, we develop a fully offline method that requires few demonstrations and captures diverse agile behaviors in a single model, simplifying the training and deployment process onto real-world robots.
Diffusion Models
Diffusion models [35, 36] are a class of generative models popularized in computer vision that have recently shown promise in robotics for imitation learning [23], offline RL [37, 38], and motion planning [23, 39]. Compared with previous generative modeling approaches such as GANs [40], diffusion models generally exhibit improved sample quality and diversity with a tradeoff in sampling time. In this work, we enhance the quality of samples generated by diffusion models while maintaining the same sampling time. This approach allows us to produce higher-quality motion samples without increasing computational cost, which is crucial for learning agile motions.
Recent works using diffusion models for robotic manipulation, such as BESO [41] and MPD2 [42], have demonstrated impressive performance and high success rates in challenging domains requiring millimeter-level precision. Our work, however, explores the additional complexity of solving dynamic tasks. Furthermore, while BESO [41] is designed for goal-conditioned tasks, the goal space in our domain is difficult to define. Consequently, our method does not rely on goal conditioning and instead adapts during sampling time.
One advantage of diffusion models is the ability to guide samples towards certain objectives at sampling time. Several prior works in robotics have shown how to add constraints for motion planning [24] and state estimation [43]. The computer vision community has also investigated using diffusion models to solve inverse problems [44, 45], where the goal is to reconstruct samples from sparse measurements. We take inspiration from these techniques used in inverse problems to formulate a novel method for guiding diffusion samples towards kinematics constraints. Our method is simple and enables the constraints to be better incorporated into the resulting diffusion samples.
Figure 2: Visualization of Diffusion Sampling Process in Air-Hockey Defend: Our KCCG method moves the conditioned points in the trajectory (shown in green) towards a predicted future puck position (shown as the orange point).
III Problem Formulation
The agile robot striking problem can be formulated as a motion planning task, where the objective is to determine a trajectory for the robot that adheres to a variety of constraints. We define $\tau=\{s_{0},s_{1},...s_{H}\}$ as a sequence of states up to a time horizon $H$. Each state $s=[q,\dot{q}]\in\mathbb{R}^{d}$ is composed of the robot’s joint positions $q$ and joint velocities $\dot{q}$. The goal of motion planning is to identify a trajectory between initial and final states that satisfy a set of costs or constraints $c_{j}(\tau)$.
Trajectory optimization-based approaches formulate the motion-planning problem as $\mathcal{J}(\tau)=\sum_{j}\lambda_{j}c_{j}(\tau)$, where the objective is to minimize the total cost $\mathcal{J}$ while satisfying the various constraints. Other approaches formulate the motion-planning problem through probabilistic inference. Here, the optimality of a state within the trajectory is represented as $\mathcal{O}_{t}$ and the goal is to sample from a posterior distribution where the $p(\tau|\mathcal{O}_{1:t})\sim p(\tau)p(\mathcal{O}_{1:t}|\tau)$, turning the motion planning problem into a conditional sampling problem [23]. We use this formulation, where the diffusion model learns the prior $p(\tau)$, and we use the guided sampling formulation [46] to conditionally sample the model on our task constraints.
Task constraints of interest for the agile striking problem include the robot kinematic constraints and ensuring contact between the end-effector and the moving object, where $b\in\mathbb{R}^{3}$ is the object’s position in Cartesian space. The forward kinematics function that maps the robot joint positions $q$ to end-effector position $x$ is denoted as $F\colon q\rightarrow x,q\in R^{d},x\in R^{3}$. Our constraint minimizes the distance between the end-effector and object positions at timestep $t$ within a time window $[t_{s},t_{e}]$ where the robot can feasibly strike the object, i.e. $c(\tau)=\min_{t\in[t_{s},t_{e}],q_{t}\in\tau}\|F(q_{t})-b_{t}\|_{2}^{2}$. Our method is amenable to other types of constraints (e.g. for obstacle avoidance).
IV Preliminaries: Score-Based Diffusion Models
Diffusion models are a class of generative models that learn to generate samples through a forward noising process and a reverse de-noising process. The forward noising process is commonly defined by: $d\tau_{t}=f(\tau_{t},t)\,dt+\sigma(t)\,dw_{t}$, where $\tau$ is the state trajectory of the robot, $f(\tau_{t},t)$ is the drift term and $\sigma(t)$ is the standard Brownian noise process. The forwards process is coupled with a backwards process: $d\tau_{t}=\left[\mu(\tau_{t},t)-\frac{1}{2}\sigma(t)^{2}\nabla_{\tau_{t}}\log p_{t}(\tau_{t})\right]dt$. In score-based diffusion models, the score function, $\nabla_{\tau_{t}}\log p_{t}(\tau_{t},t)$, is typically parametrized by a neural network $s_{\theta}(\tau,t)$. The model is trained with a score-matching loss: $\min_{\theta}\mathbb{E}_{p(\tau_{t},t)}\left[\left\|s_{\theta}(\tau_{t},t)-\nabla_{\tau_{t}}\log p(\tau_{t},t)\right\|_{2}^{2}\right]$.
The formulation provided for score-matching diffusion above is of the continuous Stochastic Differential Equation (SDE) form, where the denoising timesteps are represented in continuous time. However, the SDE can be discretized to retrieve a discrete form that matches the Denoising Diffusion Probabilistic Models (DDPM) [35] formulation commonly referred to as the Variance Preserving (VP) SDE [36]. In this formulation, the forward noising process is defined as: $\tau_{i}=\sqrt{\alpha_{i}}\tau_{i-1}+\sqrt{1-\alpha_{i}}z,z\sim\mathcal{N}(0,I)$. Following the convention from DDPM, $\beta_{i}$ represent the diagonal matrix of variances parameterizing the normal distribution, $\alpha_{i}=1-\beta_{i}$, and $\bar{\alpha}_{i}=\prod_{s=0}^{i}\alpha_{s}$. The variance parameter $\beta_{i}$ is defined by the noise variance schedule, where we use the standard cosine schedule. Complete noise is defined at $i=T$ while the data distribution is defined at $i=0$. The relationship between the score-function $s_{\theta}(\tau_{i},i)$ and the model learned through DDPM’s epsilon matching $z_{\theta}(\tau_{i},i)$ is $z_{\theta}(\tau_{i},i)=-\sqrt{1-\bar{\alpha}_{i}}s_{\theta}(\tau_{i},i)$. For brevity, we omit $i$ and write $s_{\theta}(\tau_{i})$ rather than $s_{\theta}(\tau_{i},i)$.
V Methodology
In this section, we present the technical details of our novel sampling approach (Section V-A) for constrained sampling of diffusion models.
V-A Kinematic Constraint Gradient Guidance
In this subsection, we outline the standard formulation for sampling from diffusion models without incorporating motion planning constraints. Then, we describe prior work for sampling with constraints and introduce our novel sampling method. We refer to our method as Kinematic Constraint Gradient Guidance (KCGG).
Standard Diffusion Model Sampling Procedure
First, the standard sampling procedure for the diffusion model without constraints conducts one denoising step with Equation 1.
$\tau_{i-1}=\frac{1}{\sqrt{\alpha_{i}}}\left(\tau_{i}+\left({1-\alpha_{i}}\right)s_{\theta}(\tau_{i})\right)+\sqrt{\sigma_{i}}z,z\sim\mathcal{N}(0,1)$ (1)
Here, $s_{\theta}$ denotes the score function parameterized by the diffusion model and $z$ is Gaussian noise parameterized by the denoising process.
Constraint Guided Sampling
In order to generate trajectory samples conditionally with respect to the optimality condition $p(\tau|\mathcal{O})$, prior works like Motion Planning Diffusion (MPD1) [24] and Diffuser [23] choose to directly guide the noisy trajectories $\tau_{i-1}$.
$\tau_{i-1}^{\prime}$ $=\frac{1}{\sqrt{\alpha_{i}}}\left(\tau_{i}+\left({1-\alpha_{i}}\right)s_{\theta}(\tau_{i})\right)$ (2)
$\tau_{i-1}$ $=\tau_{i-1}^{\prime}-\mathbf{\nabla_{\tau_{i-1}^{\prime}}c_{j}(\tau_{i-1}^{\prime})}+z,z\sim\mathcal{N}(0,I)$
One problem with this formulation is that the constraint projects the intermediate diffusion samples out of distribution of the diffusion model’s trained denoising paths [47]. We refer the reader to [48] for proofs on the coherence of these projection-based methods for diffusion model sampling.
Kinematic Constraint Gradient Guidance
KCGG modifies the diffusion sampling procedure to Equation 3.
$\tau_{i-1}$ $=\tau_{i-1}^{\prime}-\mathbf{\nabla_{\tau_{i}}c_{j}(\hat{\tau}_{0})}+z,z\sim\mathcal{N}(0,I)$ (3)
There are two modifications to prior work. 1) The gradient of constraints are computed for $\hat{\tau_{0}}$ rather than $\tau_{i-1}^{\prime}$ and 2) The gradient is computed with respect to $\tau_{i}$ rather than $\tau_{i-1}$.
By computing the gradient of the constraint for the predicted clean trajectory $c_{j}(\hat{\tau_{0}})$ rather than a noisy trajectory $c_{j}(\tau_{i-1}^{\prime})$, the constraints can be better informed. Crucially, the cost functions $c_{j}$ are not defined for noisy trajectories but rather noiseless trajectories. Therefore by using $\hat{\tau}_{0}$ we get a better gradient from the cost function. We can approximate $\hat{\tau_{0}}$ via $\tau_{i-1}$, as shown in Equation 4.
$\hat{\tau}_{0}\approx\frac{1}{\sqrt{\overline{\alpha_{i}}}}\left(\tau_{i}+(1-\overline{\alpha}_{i})s_{\theta}(\tau_{i},i)\right)$ (4)
Second, we compute the gradient with respect to $\tau_{i}$ rather than $\tau_{i-1}$. The trajectory $\tau_{i}$ is the noisy trajectory before the denoising step. The reason we choose $\tau_{i}$ rather than $\tau_{i-1}$ is that we derive our procedure as sampling from the posterior distribution of $p(\tau|c)$. As we define our process through a score-based approach, we are looking to compute $\nabla_{\tau_{i}}\log p(\tau|c)$. Using Bayes rule, we therefore can compute the posterior with $\nabla_{\tau_{i}}\log p(\tau_{i}|c)=\nabla_{\tau_{i}}\log p(c|\tau_{i})+\nabla_{\tau_{i}}\log p(\tau_{i})$.
Given that $\hat{\tau_{0}}$ is a function of $\tau_{i}$, our approach takes the gradient not only through the constraint function, $c_{j}$ as prior methods did but also through the diffusion model, $s_{\theta}$ itself. Intuitively, by taking the gradient through both the constraint function $c_{j}$ and the diffusion model $s_{\theta}$, we modify the entire trajectory $\tau_{i-1}$ to adhere to the constraint rather than just a portion of the trajectory. This whole-trajectory update is generally superior to updating small portions as it ensures the entire trajectory remains coherent and smooth, with all timesteps adjusted in relation to each other. It allows for global optimization of the trajectory, potentially finding better solutions that satisfy the constraint while maintaining overall motion quality. Furthermore, by updating the whole trajectory, we reduce the risk of getting stuck in local minima that may occur when only modifying small portions.
Many motion planning task constraints, such as constraining the start or ending states, only act on a few timesteps within the trajectory. Similarly, our constraints only enforce the end-effector to be at the ball at a single timestep. In traditional approaches, the function $\nabla_{\tau_{i-1}}c(\tau_{i-1})$ only updates a small portion of the predicted trajectory at each denoising timestep, where all other points receive no updates as the gradient is 0. This limited update can lead to suboptimal solutions, as it doesn’t consider the impact on the entire motion sequence. In contrast, our method’s whole-trajectory update allows for more flexible and globally optimal solutions, even when constraints are localized to specific timesteps.
Algorithm 1 Batched Sampling with KCGG
1:Constraint function $c_{j}$
2:$\boldsymbol{\tau}_{T}\leftarrow$ sample batch of trajectories from $\mathcal{N}(0,I)$
3:for all $i$ from $T$ to $1$ do
4:     $\hat{\mathbf{s}}\leftarrow s_{\theta}(\boldsymbol{\tau}_{i},i)$triangleright Predict score
5:     $\hat{\boldsymbol{\tau}}_{0}\leftarrow\frac{1}{\sqrt{\overline{\alpha}_{i}}}\left(\boldsymbol{\tau}_{i}+(1-\overline{\alpha}_{i})\hat{\mathbf{s}}\right)$triangleright Estimate $\hat{\boldsymbol{\tau}}_{0}$
6:     $\mathbf{z}\sim\mathcal{N}(0,I)$triangleright Sample Noise
7:     $\boldsymbol{\tau}^{\prime}_{i-1}\leftarrow\sqrt{\frac{\alpha_{i}(1-\overline{\alpha}_{i-1})}{1-\overline{\alpha}_{i}}}\boldsymbol{\tau}_{i}+\sqrt{\frac{\overline{\alpha}_{i-1}\beta_{i}}{1-\overline{\alpha}_{i}}}\hat{\boldsymbol{\tau}}_{0}+\tilde{\sigma}_{i}\mathbf{z}$
8:     $\boldsymbol{\tau}_{i-1}\leftarrow\boldsymbol{\tau}^{\prime}_{i-1}-\nabla_{\boldsymbol{\tau}_{i}}(c_{j}(\hat{\boldsymbol{\tau}}_{0}))$triangleright KCGG update
9:end for
10:$\tau_{0}^{*}\leftarrow\arg\min(c_{j}(\boldsymbol{\tau}_{0}))$triangleright Batch filter
11:return $\tau_{0}^{*}$
V-B Sampling Procedure
We describe our guidance procedure in Algorithm 1. We predict a batch of trajectories $\hat{\tau}_{0}$ with our network $s_{\theta}(\tau_{i},i)$ (Lines 4 - 5). An intermediate sample $\tau_{i-1}^{\prime}$ is computed using the standard DDPM procedure (Lines 6 - 7). Then, the sample is updated by KCGG in Line 8. The function $F$ represents the forward kinematics chain for the robot arm. We define our constraint function as $c(\tau)=\min_{t\in[t_{s},t_{e}],q_{t}\in\tau}\|F(q_{t})-b_{t}\|_{2}^{2}$, which minimizes the difference between the robot end-effector and ball prediction location at certain points in time defined by the window, $[t_{s},t_{e}]$. Finally, we select a single best trajectory $\tau_{0}^{*}$ based on the computed norm $N$ from the final denoising step (Line 10). This allows us to utilize the computational efficiency of batched operations on the GPU and select the trajectory that best matches the given end-effector constraint. We refer to this step as “Batch Filter” and ablate this feature.
VI Results and Discussion
In this section, we seek to address two primary questions: 1) Do diffusion models have the capacity to generate diverse behaviors for executing the same task in multiple ways? 2) Can KCGG facilitate trajectory generation from diffusion models to improve agile robotic tasks?
VI-A Agile Robotic Domains
To evaluate our method, we perform experiments on two distinct domains: AirHockey and Table-Tennis.
VI-A1 Simulated Domain (AirHockey)
We evaluate the air hockey simulation domain [6] on two tasks, Hit and Defend. In Hit, the objective is to score a goal with the desired shot type, Straight or Angled. If a goal is scored with the shot type behavior, it is considered a Success. In Defend, the goal is to block an incoming puck. Block is defined as when the robot’s end-effector successfully makes contact with the puck. Models are trained on 200 (100 per shot type) demonstrations in Hit and 100 demonstrations in Defend.
Method Straight Angle Combined
ProMP Unconditioned - - 8.1%
Behavioral Cloning (BC) [49] 9.5% 0.8% 5.2%
ACT [19] 89.4% 65.2% 77.3%
Conditional Diffusion Model 98.3% 82.4% 90.4%
TABLE I: AirHockey Hit Results: Comparison of Success rates on Straight and Angled shot types.
VI-A2 Real-World Domain (Table Tennis)
A 7-DOF Barrett WAM system from [30] with perception improvements [50] is used to return table tennis serves from a ball launcher. We collect diverse training data through kinesthetic teaching and record four types of shots: Push, Lob, Slice, and Topspin. Balls are launched to two striking locations on the table: Center and Left. Models are trained on 120 demonstrations (15 per condition) and evaluated on three metrics: Success, Hit, and Correct Stroke rate. Success is defined as a ball return within the spin and height distribution of the demonstrated shot type’s data. A Hit is recorded when the ball contacts the end-effector, but fails the Success criteria. Finally, Correct Stroke is a sampled stroke that matches the desired shot type’s joint trajectory distribution and striking location.
VI-B Baselines
We benchmarked against the following LfD baselines: Behavioral cloning (BC) policies with Gaussian Mixture Models (GMM) [49], Probabilistic Movement Primitives (ProMP) [26], and Action Chunking with Transformers (ACT) [19]. We also compare against five diffusion baselines: Diffuser [23], Motion Planning Diffusion (MPD1) [24], Diffusion Policy [51], BEbhavior generation with ScOre-based Diffusion Policies (BESO) [41] and Movement Primitive Diffusion (MPD2) [42].
Method Block Rate $\uparrow$ Time (s) T ms/step
Imitation Learning Baselines
Behavioral Cloning (BC) [49] 15.5% - - -
ACT [19] 79.4% - - -
Unconditioned Diffusion Models
Unconstrained w/o Batch Filter 12.9% 0.159 20 7.95
Unconstrained 62.6% 0.161 20 8.05
MPD1 [24] 63.5% 0.190 16 11.88
KCGG (Ours) 85.2% 0.189 10 18.90
Puck Conditioned Diffusion Models
Unconstrained w/o Batch Filter 46.2% 0.162 20 8.10
Unconstrained 46.6% 0.163 20 8.15
MPD1 [24] 58.2% 0.191 16 11.94
BESO [41] 57.4% 0.191 35 5.45
MPD2 [42] 66.3% 0.192 21 9.14
KCGG (Ours) 47.3% 0.188 10 18.80
TABLE II: Air-Hockey Defend Results: We compare our constraint sampling method against imitation learning and diffusion baselines. “T” is the number of diffusion timesteps.
VI-C Simulation Results: AirHockey
In this section, we discuss results on the simulated AirHockey Hit and Defend tasks. For each task, we test on 1000 initial start configurations. Our results for Hit (Table I) show that the diffusion model captures diverse motions best, achieving highest success rates for each shot type. Results for the Defend task (Table II), show KCGG enables unconditioned diffusion models to outperform all baselines, increasing hit rate by 18.9% over the best baseline. We use a fixed wall-clock time budget of 200 ms and tune sampling steps accordingly. Interestingly, KCGG does not improve puck-conditioned models. We hypothesize the conditioned base model overfits the score function to puck observations.
Figure 3: Comparison of Planned Trajectory: KCGG maintains a smooth joint angle profile in the constrained region (dotted lines) while constraint guidance introduces discontinuities.
Figure 3 shows a comparison of the joint angle trajectories generated by KCGG and the original constraint guidance, revealing KCGG more effectively achieves the desired end-effector trajectory while adhering to the data distribution. The trajectory produced by the original constraint guidance exhibits significant inconsistencies in the constrained region whereas KCGG samples a smooth trajectory.
VI-D Sampling Time Ablation
We perform a sensitivity analysis comparing against other diffusion baselines by varying the number of diffusion timesteps (Figure 4). Despite requiring more time per step, KCGG outperforms diffusion baselines across all sampling time budgets, except for MPD2 with sampling time budget less than $70$ ms. Wall-clock time is evaluated as sampling step time vary for each method. All experiments were run on a single Nvidia GTX 1080.
Figure 4: Sampling Speed vs Performance: We compare the Block Rate in Air-Hockey Defend across sampling times. KCGG scales to better performance with sampling time.
As the number of diffusion timesteps increases, the performance of the baseline methods plateau. In contrast, KCGG scales effectively, reaching a block rate nearing 85%, while a method such as MPD2 does not, despite performing better with few sampling steps. Additional denoising steps allows KCGG to better adhere the trajectory samples to both the training data distribution and task constraints.
Left Push Left Lob Left Slice Left Topspin
Method H S CS H S CS H S CS H S CS
ProMP [26] 50% 20% - 80% 30% - 42% 26% - 66% 40% -
ACT [19] 0% 0% 66% 64% 40% 88% 82% 14% 100% 70% 24% 88%
Unconstrained Diffusion 28% 8% 72% 38% 6% 46% 10% 4% 12% 14% 10% 22%
MPD1 [24] 36% 22% 100% 70% 16% 100% 78% 8% 100% 52% 16% 100%
BESO [41] 0% 0% 0% 0% 0% 0% 0% 0% 2% 0% 0% 0%
BESO w/ ball oracle [41] 60% 46% 100% 62% 28% 100% 74% 2% 100% 44% 10% 100%
MPD2 [42] 72% 42% 94% 2% 2% 2% 4% 0% 6% 2% 2% 2%
MPD2 w/ ball oracle [42] 46% 20% 100% 80% 32% 100% 52% 0% 100% 56% 34% 100%
KCGG (Ours) 76% 50% 100% 96% 60% 100% 64% 26% 100% 52% 26% 100%
Center Push Center Lob Center Slice Center Topspin
Method H S CS H S CS H S CS H S CS
ProMP [26] 84% 18% - 58% 24% - 76% 0% - 74% 2% -
ACT [19] 56% 10% 90% 70% 26% 74% 28% 0% 34% 40% 22% 62%
Unconstrained Diffusion 42% 6% 42% 36% 10% 40% 38% 0% 98% 50% 16% 82%
MPD1 [24] 66% 0% 94% 92% 18% 100% 84% 0% 100% 72% 0% 100%
BESO [41] 2% 0% 2% 10% 4% 92% 0% 0% 0% 0% 0% 0%
BESO w/ ball oracle [41] 94% 0% 100% 92% 20% 100% 40% 0% 100% 44% 18% 100%
MPD2 [42] 2% 0% 4% 44% 14% 86% 80% 0% 100% 2% 0% 2%
MPD2 w/ ball oracle [42] 66% 0% 100% 64% 12% 100% 30% 0% 100% 74% 10% 100%
KCGG (Ours) 100% 10% 100% 100% 32% 100% 72% 0% 100% 70% 22% 94%
TABLE III: Table Tennis Results: We compare the hit rates (H), success rates (S), and correct stroke rates (CS) for four different shot types across two different ball launcher configurations (left and center launches).
VI-E Real-world Results: Table Tennis
We report results comparing diffusion models to imitation learning baselines in Table III with 50 samples per task. Expert demonstrations are used instead of traditional heuristics, which lack the flexibility to generate diverse table-tennis stroke profiles and are hard to design. Based on Air-Hockey Defend findings, we conditioned our models only on stroke type to avoid over-parameterizing the score function.
In terms of task performance, KCGG shows an average 17.3% improvement in success rate across all tasks, outperforming or matching the highest baseline success rates in six of eight tasks. Specifically, KCGG surpasses prior diffusion constraint guidance methods (MPD1) with a notable 124.5% improvement. We also note a modest average 4% improvement in hit rate over the baselines, attributed mainly to high hit rates in the push and lob strokes.
KCGG excels in most strokes but has a lower hit rate in slice and topspin. This may be due to tighter timing requirements in these strokes, where less overlap between the ball and end-effector trajectories is present, posing challenges for KCGG in generating out-of-distribution trajectories (see Figure 3). While prior methods like MPD1 can adjust trajectories to meet hit points, they do not necessarily enhance success rates as they alter the stroke profile.
We evaluated KCGG’s ability to express multimodal behaviors across four stroke types and two ball launch configurations. KCGG consistently demonstrated the desired behavior in each task, shown by high correct stroke rates. Conversely, ACT, despite conditioning on stroke style, struggled to generate correct left/center strokes from ball observations, yielding a lower correct stroke rate. For BESO and MPD2, we observed that conditioning on ball observation history caused overfitting and collapse into certain stroke modalities (e.g. BESO mostly samples Center Lob strokes), illustrating the importance of constraint-guidance. We created stronger variants of these baselines by replacing ball history conditioning with ”ball oracle” conditioning, a categorical label indicating Left or Center launch distribution from the demonstration. This provides direct information on the ball launch location, which significantly improved BESO and MPD2 performance, removing mode collapse for a stronger comparison. However, KCGG still outperformed these variants on average in hit and success rate. ProMP, which was not designed to capture multimodal behavior, was trained on data specific to each task, thus we exclude it from correct stroke rate evaluations. KCGG, however, effectively modeled trajectories across all conditions, accurately selecting the appropriate stroke based on the constraints.
VII Limitations and Future Work
The primary drawback of our method is that we cannot generate trajectories that deviate significantly from the training data distribution. Because our approach relies on computing gradients through the trained score function, it produces samples that closely resemble behaviors in the training dataset. While this property ensures that the generated motions remain within the demonstrated range, it restricts the robot’s adaptability to unseen scenarios. Future work could address this by using KCGG to warm-start RL.
VIII Conclusion
Our work demonstrates KCGG’s effectiveness on small expert datasets in agile robotic tasks. Our algorithm is validated in both a simulated air hockey domain and on a real table tennis robot. Our real robot experiments show that we can learn diverse policies without access to a high-fidelity simulator. These contributions address critical issues such as sample inefficiency and performance in dynamic settings, promising broader applications in real-world robotics.

### v2

@@
 We demonstrate the effectiveness of our approach for time-critical robotic tasks by evaluating KCGG in two challenging domains: simulated air hockey and real table tennis.
-In simulated air hockey, we achieved a 25.4% increase in block rate, while in table tennis, we saw a 17.3% increase in success rate compared to imitation learning baselines.
+In simulated air hockey, we achieved a 25.4% increase in block rate, while in table tennis, we achieved a 17.3% increase in success rate compared to imitation learning baselines.
 I Introduction
@@
 Deep learning-based LfD algorithms also typically do not have a formulation to incorporate novel kinematic or dynamic constraints at test time.
-Figure 1: Overview of KCGG: Our diffusion model was trained on four kinesthetically demonstrated strokes.
-At inference, for each denoising step, our KCGG method computes the cost given the constraint function and intermediate trajectory.
-Then, the gradient with respect to the prior trajectory through the diffusion model is computed to update the sample.
+Figure 1: Overview of KCGG: At inference, KCGG computes the cost given the constraint and intermediate trajectory at each denoising step.
+The gradient with respect to the prior trajectory through the diffusion model is used to update the sample.
 In this paper, we develop a novel imitation learning based approach that enables robots to learn from small offline datasets to perform agile tasks in dynamical environments.
-We ground our work in application to striking motions evaluated in established challenge domains for robotics: a virtual air hockey domain [20, 9] and a physical table tennis domain [21, 22].
+We ground our work in an application to striking motions evaluated in established challenging domains for robotics: a virtual air hockey domain [20, 9] and a physical table tennis domain [21, 22].
 We show that our diffusion-based approach can learn to condition trajectories to meet environment constraints and perform tasks difficult to specify (e.g.
@@
   • We show KCGG improves a real-world table tennis robot’s ability to return ping-pong balls.
-Our method improves success rate by 124.5% over diffusion baselines [23, 24] and 17.3% over state of the art imitation learning baselines [19], demonstrating efficacy in agile robotic tasks.
+Our method improves success rate by 124.5% over diffusion baselines [23, 24] and 17.3% over state-of-the-art imitation learning baselines [19], demonstrating efficacy in agile robotic tasks.
 II Related Work
@@
 MP-based LfD has been explored extensively for agile striking motions in table tennis by learning movement primitives.
-For example, Mixture of Movement Primitives (MoMP) [22, 25] learns to select a dynamic movement primitive (DMP) from a library of pre-trained DMPs to generate joint trajectories for returning the ball.
+For example, Mixture of Movement Primitives (MoMP) [22, 25] learns to select a dynamic movement primitive (DMP) from a library of pre-trained DMPs to generate joint trajectories for ball returns.
 Probabilistic movement primitives (ProMP) [26], which model the variability in movements as trajectory distributions, were adapted to generate striking motions that maximize the likelihood of ball return [27, 28, 29, 30].
@@
 lob in table tennis) is difficult without a hierarchical approach.
-Recently, RL methods that are utilizing heavy computational resources are enabling systems to explore in task space without using demonstrated expert knowledge [21, 31, 32, 33, 34].
-However, most of them require large training datasets [21, 32], carefully shaped reward [21], or human behavior modeling [34] which require a simulation.
+Recently, RL methods are enabling systems to explore in task space without using demonstrated expert knowledge [21, 31, 32, 33, 34].
+However, most of them require large training datasets [21, 32], carefully shaped rewards [21], or human behavior modeling [34] which require a simulation.
 In contrast, we develop a fully offline method that requires few demonstrations and captures diverse agile behaviors in a single model, simplifying the training and deployment process onto real-world robots.
 Diffusion Models
-Diffusion models [35, 36] are a class of generative models popularized in computer vision that have recently shown promise in robotics for imitation learning [23], offline RL [37, 38], and motion planning [23, 39].
+Diffusion models [35, 36] are a class of generative models that have recently shown promise in robotics for imitation learning [23], offline RL [37, 38], and motion planning [23, 39].
 Compared with previous generative modeling approaches such as GANs [40], diffusion models generally exhibit improved sample quality and diversity with a tradeoff in sampling time.
-In this work, we enhance the quality of samples generated by diffusion models while maintaining the same sampling time.
-This approach allows us to produce higher-quality motion samples without increasing computational cost, which is crucial for learning agile motions.
-Recent works using diffusion models for robotic manipulation, such as BESO [41] and MPD2 [42], have demonstrated impressive performance and high success rates in challenging domains requiring millimeter-level precision.
+In this work, we enhance the quality of samples generated by diffusion models while maintaining similar sampling time.
+This approach produces higher-quality samples without increasing computational cost, which is crucial for learning agile motions.
+Recent works using diffusion models for robotic manipulation, such as BESO [41] and MPD2 [42], demonstrated impressive performance and high success rates in challenging domains requiring millimeter-level precision.
 Our work, however, explores the additional complexity of solving dynamic tasks.
@@
 The computer vision community has also investigated using diffusion models to solve inverse problems [44, 45], where the goal is to reconstruct samples from sparse measurements.
-We take inspiration from these techniques used in inverse problems to formulate a novel method for guiding diffusion samples towards kinematics constraints.
+We take inspiration from these techniques used in inverse problems to formulate a novel method for guiding diffusion samples towards kinematics constraints and evaluate this method in a difficult agile robotics domain.
 Our method is simple and enables the constraints to be better incorporated into the resulting diffusion samples.
-Figure 2: Visualization of Diffusion Sampling Process in Air-Hockey Defend: Our KCCG method moves the conditioned points in the trajectory (shown in green) towards a predicted future puck position (shown as the orange point).
+Figure 2: Visualization of Diffusion Sampling Process in Air-Hockey Defend: KCGG gradually shifts the conditioned points in the trajectory (shown in green) towards a predicted future puck position (shown as the orange point).
 III Problem Formulation
 The agile robot striking problem can be formulated as a motion planning task, where the objective is to determine a trajectory for the robot that adheres to a variety of constraints.
-We define $\tau=\{s_{0},s_{1},...s_{H}\}$ as a sequence of states up to a time horizon $H$.
-Each state $s=[q,\dot{q}]\in\mathbb{R}^{d}$ is composed of the robot’s joint positions $q$ and joint velocities $\dot{q}$.
-The goal of motion planning is to identify a trajectory between initial and final states that satisfy a set of costs or constraints $c_{j}(\tau)$.
-Trajectory optimization-based approaches formulate the motion-planning problem as $\mathcal{J}(\tau)=\sum_{j}\lambda_{j}c_{j}(\tau)$, where the objective is to minimize the total cost $\mathcal{J}$ while satisfying the various constraints.
+We define $\tau=\{s_{0},s_{1},...s_{H}\}$ as a sequence of states up to a time horizon, $H$.
+Each state $s=[q,\dot{q}]\in\mathbb{R}^{d}$ is composed of the robot’s joint positions, $q$, and joint velocities, $\dot{q}$.
+The goal of motion planning is to identify a trajectory between initial and final states that satisfy a set of costs or constraints, $c_{j}(\tau)$.
+The objective is to minimize the total cost $\mathcal{J}(\tau)=\sum_{j}\lambda_{j}c_{j}(\tau)$, subject to kinematic or other constraints.
 Other approaches formulate the motion-planning problem through probabilistic inference.
-Here, the optimality of a state within the trajectory is represented as $\mathcal{O}_{t}$ and the goal is to sample from a posterior distribution where the $p(\tau|\mathcal{O}_{1:t})\sim p(\tau)p(\mathcal{O}_{1:t}|\tau)$, turning the motion planning problem into a conditional sampling problem [23].
-We use this formulation, where the diffusion model learns the prior $p(\tau)$, and we use the guided sampling formulation [46] to conditionally sample the model on our task constraints.
-Task constraints of interest for the agile striking problem include the robot kinematic constraints and ensuring contact between the end-effector and the moving object, where $b\in\mathbb{R}^{3}$ is the object’s position in Cartesian space.
-The forward kinematics function that maps the robot joint positions $q$ to end-effector position $x$ is denoted as $F\colon q\rightarrow x,q\in R^{d},x\in R^{3}$.
-Our constraint minimizes the distance between the end-effector and object positions at timestep $t$ within a time window $[t_{s},t_{e}]$ where the robot can feasibly strike the object, i.e.
+Here, the optimality of a state within the trajectory is represented as $\mathcal{O}_{t}$ and the goal is to sample from a posterior distribution, $p(\tau|\mathcal{O}_{1:t})\sim p(\tau)p(\mathcal{O}_{1:t}|\tau)$ [23].
+We use this formulation, where the diffusion model learns the prior, $p(\tau)$, and we use the guided sampling formulation [46] to conditionally sample the model on our task constraints.
+A task constraint of interest for the agile striking problem is to ensure contact between the end-effector and the moving object, where $b\in\mathbb{R}^{3}$ is the object’s position in Cartesian space.
+The forward kinematics function that maps the robot joint positions, $q$, to end-effector position, $x$, is denoted as $F\colon q\rightarrow x,q\in R^{d},x\in R^{3}$.
+Our constraint minimizes the distance between the end-effector and object positions at a timestep, $t$, within a time window, $[t_{s},t_{e}]$, where the robot can feasibly strike the object, i.e.
 $c(\tau)=\min_{t\in[t_{s},t_{e}],q_{t}\in\tau}\|F(q_{t})-b_{t}\|_{2}^{2}$.
@@
 The forwards process is coupled with a backwards process: $d\tau_{t}=\left[\mu(\tau_{t},t)-\frac{1}{2}\sigma(t)^{2}\nabla_{\tau_{t}}\log p_{t}(\tau_{t})\right]dt$.
-In score-based diffusion models, the score function, $\nabla_{\tau_{t}}\log p_{t}(\tau_{t},t)$, is typically parametrized by a neural network $s_{\theta}(\tau,t)$.
+In score-based diffusion models, the score function, $\nabla_{\tau_{t}}\log p_{t}(\tau_{t},t)$, is typically parametrized by a neural network, $s_{\theta}(\tau,t)$.
 The model is trained with a score-matching loss: $\min_{\theta}\mathbb{E}_{p(\tau_{t},t)}\left[\left\|s_{\theta}(\tau_{t},t)-\nabla_{\tau_{t}}\log p(\tau_{t},t)\right\|_{2}^{2}\right]$.
@@
 Following the convention from DDPM, $\beta_{i}$ represent the diagonal matrix of variances parameterizing the normal distribution, $\alpha_{i}=1-\beta_{i}$, and $\bar{\alpha}_{i}=\prod_{s=0}^{i}\alpha_{s}$.
-The variance parameter $\beta_{i}$ is defined by the noise variance schedule, where we use the standard cosine schedule.
+The variance parameter, $\beta_{i}$, is defined by the noise variance schedule, where we use the standard cosine schedule.
 Complete noise is defined at $i=T$ while the data distribution is defined at $i=0$.
-The relationship between the score-function $s_{\theta}(\tau_{i},i)$ and the model learned through DDPM’s epsilon matching $z_{\theta}(\tau_{i},i)$ is $z_{\theta}(\tau_{i},i)=-\sqrt{1-\bar{\alpha}_{i}}s_{\theta}(\tau_{i},i)$.
-For brevity, we omit $i$ and write $s_{\theta}(\tau_{i})$ rather than $s_{\theta}(\tau_{i},i)$.
+The relationship between the score-function, $s_{\theta}(\tau_{i},i)$, and the model learned through DDPM’s epsilon matching, $z_{\theta}(\tau_{i},i)$, is $z_{\theta}(\tau_{i},i)=-\sqrt{1-\bar{\alpha}_{i}}s_{\theta}(\tau_{i},i)$.
+For brevity, we omit $i$ and denote $s_{\theta}(\tau_{i})$ as $s_{\theta}(\tau_{i},i)$.
 V Methodology
@@
 In this subsection, we outline the standard formulation for sampling from diffusion models without incorporating motion planning constraints.
-Then, we describe prior work for sampling with constraints and introduce our novel sampling method.
-We refer to our method as Kinematic Constraint Gradient Guidance (KCGG).
+Then, we describe prior work for sampling with constraints and introduce our novel sampling approach.
 Standard Diffusion Model Sampling Procedure
-First, the standard sampling procedure for the diffusion model without constraints conducts one denoising step with Equation 1.
-$\tau_{i-1}=\frac{1}{\sqrt{\alpha_{i}}}\left(\tau_{i}+\left({1-\alpha_{i}}\right)s_{\theta}(\tau_{i})\right)+\sqrt{\sigma_{i}}z,z\sim\mathcal{N}(0,1)$ (1)
-Here, $s_{\theta}$ denotes the score function parameterized by the diffusion model and $z$ is Gaussian noise parameterized by the denoising process.
+Each denoising step in a standard, unconstrained diffusion model follows Eq.
+1, where $s_{\theta}$ denotes the score function parameterized by the diffusion model and $z\sim\mathcal{N}(0,1)$ is Gaussian noise parameterized by the denoising process.
+$\tau_{i-1}=\frac{1}{\sqrt{\alpha_{i}}}\left(\tau_{i}+\left({1-\alpha_{i}}\right)s_{\theta}(\tau_{i})\right)+\sqrt{\sigma_{i}}z$ (1)
 Constraint Guided Sampling
-In order to generate trajectory samples conditionally with respect to the optimality condition $p(\tau|\mathcal{O})$, prior works like Motion Planning Diffusion (MPD1) [24] and Diffuser [23] choose to directly guide the noisy trajectories $\tau_{i-1}$.
-$\tau_{i-1}^{\prime}$ $=\frac{1}{\sqrt{\alpha_{i}}}\left(\tau_{i}+\left({1-\alpha_{i}}\right)s_{\theta}(\tau_{i})\right)$ (2)
-$\tau_{i-1}$ $=\tau_{i-1}^{\prime}-\mathbf{\nabla_{\tau_{i-1}^{\prime}}c_{j}(\tau_{i-1}^{\prime})}+z,z\sim\mathcal{N}(0,I)$
-One problem with this formulation is that the constraint projects the intermediate diffusion samples out of distribution of the diffusion model’s trained denoising paths [47].
+To generate trajectory samples conditionally with respect to the optimality condition, $p(\tau|\mathcal{O})$, prior works choose to directly guide the noisy trajectories, $\tau_{i-1}$, as follows in Eq.
+2.
+$\tau_{i-1}^{\prime}$ $=\frac{1}{\sqrt{\alpha_{i}}}\left(\tau_{i}+\left({1-\alpha_{i}}\right)s_{\theta}(\tau_{i})\right),$ (2)
+$\tau_{i-1}$ $=\tau_{i-1}^{\prime}-\mathbf{\nabla_{\tau_{i-1}^{\prime}}c_{j}(\tau_{i-1}^{\prime})}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sqrt{\sigma_{i}}}z$
+However, with this formulation, the constraint projects the intermediate diffusion samples out of distribution from the diffusion model’s trained denoising paths [47].
 We refer the reader to [48] for proofs on the coherence of these projection-based methods for diffusion model sampling.
 Kinematic Constraint Gradient Guidance
-KCGG modifies the diffusion sampling procedure to Equation 3.
-$\tau_{i-1}$ $=\tau_{i-1}^{\prime}-\mathbf{\nabla_{\tau_{i}}c_{j}(\hat{\tau}_{0})}+z,z\sim\mathcal{N}(0,I)$ (3)
+KCGG modifies the diffusion sampling procedure to Eq. 3.
+$\tau_{i-1}$ $=\tau_{i-1}^{\prime}-\mathbf{\nabla_{\tau_{i}}c_{j}(\hat{\tau}_{0})}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sqrt{\sigma_{i}}}z$ (3)
 There are two modifications to prior work.
 1) The gradient of constraints are computed for $\hat{\tau_{0}}$ rather than $\tau_{i-1}^{\prime}$ and 2) The gradient is computed with respect to $\tau_{i}$ rather than $\tau_{i-1}$.
-By computing the gradient of the constraint for the predicted clean trajectory $c_{j}(\hat{\tau_{0}})$ rather than a noisy trajectory $c_{j}(\tau_{i-1}^{\prime})$, the constraints can be better informed.
-Crucially, the cost functions $c_{j}$ are not defined for noisy trajectories but rather noiseless trajectories.
+By computing the gradient of the constraint for the predicted clean trajectory, $c_{j}(\hat{\tau_{0}})$, rather than a noisy trajectory, $c_{j}(\tau_{i-1}^{\prime})$, the constraints can be better informed.
+Crucially, the cost functions, $c_{j}$, are not defined for noisy trajectories but rather noiseless trajectories.
 Therefore by using $\hat{\tau}_{0}$ we get a better gradient from the cost function.
@@
 Second, we compute the gradient with respect to $\tau_{i}$ rather than $\tau_{i-1}$.
-The trajectory $\tau_{i}$ is the noisy trajectory before the denoising step.
-The reason we choose $\tau_{i}$ rather than $\tau_{i-1}$ is that we derive our procedure as sampling from the posterior distribution of $p(\tau|c)$.
-As we define our process through a score-based approach, we are looking to compute $\nabla_{\tau_{i}}\log p(\tau|c)$.
+$\tau_{i}$ is the noisy trajectory before the denoising step.
+We choose $\tau_{i}$ rather than $\tau_{i-1}$ because we derive our procedure as sampling from the posterior distribution of $p(\tau|c)$.
+As we define our process through a score-based approach, we compute $\nabla_{\tau_{i}}\log p(\tau|c)$.
 Using Bayes rule, we therefore can compute the posterior with $\nabla_{\tau_{i}}\log p(\tau_{i}|c)=\nabla_{\tau_{i}}\log p(c|\tau_{i})+\nabla_{\tau_{i}}\log p(\tau_{i})$.
-Given that $\hat{\tau_{0}}$ is a function of $\tau_{i}$, our approach takes the gradient not only through the constraint function, $c_{j}$ as prior methods did but also through the diffusion model, $s_{\theta}$ itself.
-Intuitively, by taking the gradient through both the constraint function $c_{j}$ and the diffusion model $s_{\theta}$, we modify the entire trajectory $\tau_{i-1}$ to adhere to the constraint rather than just a portion of the trajectory.
-This whole-trajectory update is generally superior to updating small portions as it ensures the entire trajectory remains coherent and smooth, with all timesteps adjusted in relation to each other.
-It allows for global optimization of the trajectory, potentially finding better solutions that satisfy the constraint while maintaining overall motion quality.
-Furthermore, by updating the whole trajectory, we reduce the risk of getting stuck in local minima that may occur when only modifying small portions.
-Many motion planning task constraints, such as constraining the start or ending states, only act on a few timesteps within the trajectory.
-Similarly, our constraints only enforce the end-effector to be at the ball at a single timestep.
-In traditional approaches, the function $\nabla_{\tau_{i-1}}c(\tau_{i-1})$ only updates a small portion of the predicted trajectory at each denoising timestep, where all other points receive no updates as the gradient is 0.
-This limited update can lead to suboptimal solutions, as it doesn’t consider the impact on the entire motion sequence.
-In contrast, our method’s whole-trajectory update allows for more flexible and globally optimal solutions, even when constraints are localized to specific timesteps.
 Algorithm 1 Batched Sampling with KCGG
@@
 11:return $\tau_{0}^{*}$
+Given that $\hat{\tau_{0}}$ is a function of $\tau_{i}$, our approach takes the gradient not only through the constraint function, $c_{j}$, as prior methods did, but also through the diffusion model, $s_{\theta}$.
+Intuitively, by taking the gradient through both $c_{j}$ and $s_{\theta}$, the entire trajectory, $\tau_{i-1}$, adheres to the constraint rather than just a portion of the trajectory.
+This whole-trajectory update is generally superior to updating small portions as it ensures the entire trajectory remains coherent and smooth.
+Furthermore, by updating the whole trajectory, we reduce the risk of getting stuck in local minima that may occur when only modifying small portions.
+Many motion planning task constraints, e.g.
+constraints on goal states, only act on a subset of the trajectory.
+In prior approaches, the function, $\nabla_{\tau_{i-1}}c(\tau_{i-1})$ only updates the constrained subset of the predicted trajectory at each denoising step.
+However, this can lead to suboptimal solutions, as it does not consider the impact on the entire motion sequence.
+In contrast, our method’s whole-trajectory update allows for more flexible and globally optimal solutions, even when constraints are localized to specific timesteps.
 V-B Sampling Procedure
 We describe our guidance procedure in Algorithm 1.
-We predict a batch of trajectories $\hat{\tau}_{0}$ with our network $s_{\theta}(\tau_{i},i)$ (Lines 4 - 5).
-An intermediate sample $\tau_{i-1}^{\prime}$ is computed using the standard DDPM procedure (Lines 6 - 7).
+We predict a batch of trajectories, $\hat{\tau}_{0}$, with our network, $s_{\theta}(\tau_{i},i)$ (Lines 4 - 5).
+An intermediate sample, $\tau_{i-1}^{\prime}$, is computed using the standard DDPM procedure (Lines 6 - 7).
 Then, the sample is updated by KCGG in Line 8.
@@
 We define our constraint function as $c(\tau)=\min_{t\in[t_{s},t_{e}],q_{t}\in\tau}\|F(q_{t})-b_{t}\|_{2}^{2}$, which minimizes the difference between the robot end-effector and ball prediction location at certain points in time defined by the window, $[t_{s},t_{e}]$.
-Finally, we select a single best trajectory $\tau_{0}^{*}$ based on the computed norm $N$ from the final denoising step (Line 10).
+Finally, we select a single best trajectory, $\tau_{0}^{*}$, based on the computed norm, $N$, from the final denoising step (Line 10).
 This allows us to utilize the computational efficiency of batched operations on the GPU and select the trajectory that best matches the given end-effector constraint.
 We refer to this step as “Batch Filter” and ablate this feature.
+Hyperparameter Value
+Learning Rate 2e-4
+Planning Horizon 100
+Batch Size 32
+Number of Training Diffusion Steps 100
+Total Training Steps 10e4
+Optimizer Adam
+U-Net Convolutional Blocks (1, 4, 8)
+TABLE I: Architecture Hyperparameters in both domains
 VI Results and Discussion
@@
 VI-A Agile Robotic Domains
-To evaluate our method, we perform experiments on two distinct domains: AirHockey and Table-Tennis.
+We evaluate on two domains: AirHockey and Table Tennis.
+In each domain, the object pose is estimated by a Kalman Filter [49].
+A predicted path of the object is obtained from a linear dynamics rollout in AirHockey and a ball dynamics model [50] in Table Tennis.
+We form the time window, $[t_{s},t_{e}]$, around a predetermined hit plane derived from the mean location of object hits in the collected demonstrations.
 VI-A1 Simulated Domain (AirHockey)
@@
 In Hit, the objective is to score a goal with the desired shot type, Straight or Angled.
-If a goal is scored with the shot type behavior, it is considered a Success.
-In Defend, the goal is to block an incoming puck.
-Block is defined as when the robot’s end-effector successfully makes contact with the puck.
+A goal scored with the desired shot type is considered a Success.
+In Defend, the goal is to block an incoming puck that is initialized with random position and velocity.
+A Block is defined as when the robot’s end-effector successfully makes contact with the puck.
 Models are trained on 200 (100 per shot type) demonstrations in Hit and 100 demonstrations in Defend.
+Demonstrations are collected from a heuristic motion planner.
+VI-A2 Real-World Domain (Table Tennis)
+A 7-DOF Barrett WAM system from [30] with perception improvements [51] is used to return serves from a ball launcher.
+Training data is collected through kinesthetic teaching, with four types of shots: Push, Lob, Slice, and Topspin.
+Balls are launched to two locations on the table: Center and Left, with natural ball variance from the launcher.
+Models are trained on 120 demonstrations (15 per condition) and evaluated on three metrics: Success, Hit, and Correct Stroke rate.
+Success is defined as a ball return within the spin and height distribution of the shot type’s demonstrations.
+A Hit is recorded when the ball contacts the end-effector, but fails the Success criteria.
+Finally, Correct Stroke is a sample that matches the desired shot type’s joint trajectory distribution and striking location.
 Method Straight Angle Combined
 ProMP Unconditioned - - 8.1%
-Behavioral Cloning (BC) [49] 9.5% 0.8% 5.2%
+Behavioral Cloning (BC-GMM) [52] 9.5% 0.8% 5.2%
 ACT [19] 89.4% 65.2% 77.3%
-Conditional Diffusion Model 98.3% 82.4% 90.4%
-TABLE I: AirHockey Hit Results: Comparison of Success rates on Straight and Angled shot types.
-VI-A2 Real-World Domain (Table Tennis)
-A 7-DOF Barrett WAM system from [30] with perception improvements [50] is used to return table tennis serves from a ball launcher.
-We collect diverse training data through kinesthetic teaching and record four types of shots: Push, Lob, Slice, and Topspin.
-Balls are launched to two striking locations on the table: Center and Left.
-Models are trained on 120 demonstrations (15 per condition) and evaluated on three metrics: Success, Hit, and Correct Stroke rate.
-Success is defined as a ball return within the spin and height distribution of the demonstrated shot type’s data.
-A Hit is recorded when the ball contacts the end-effector, but fails the Success criteria.
-Finally, Correct Stroke is a sampled stroke that matches the desired shot type’s joint trajectory distribution and striking location.
-VI-B Baselines
-We benchmarked against the following LfD baselines: Behavioral cloning (BC) policies with Gaussian Mixture Models (GMM) [49], Probabilistic Movement Primitives (ProMP) [26], and Action Chunking with Transformers (ACT) [19].
-We also compare against five diffusion baselines: Diffuser [23], Motion Planning Diffusion (MPD1) [24], Diffusion Policy [51], BEbhavior generation with ScOre-based Diffusion Policies (BESO) [41] and Movement Primitive Diffusion (MPD2) [42].
+Diffusion (Base) 98.3% 82.4% 90.4%
+TABLE II: AirHockey Hit Results: Comparison of Success rates on Straight and Angled shot types.
+VI-B Experiment Details and Baselines
+Due to the tight timing requirements of our domains, we use a mixture of both [23] and [53] as our baseline diffusion model, which we refer to as “Diffusion (Base)”.
+We utilize the 1D Convolutional Temporal U-Net from [23] due to its faster sampling speed over attention-based models.
+For conditioned models, we incorporate Feature-Wise Linear Modulation (FiLM) Layers [54] to modulate the convolutional features within the U-Net, as suggested by [53].
+Conditioned models use state-based observations of joint positions, stroke type, and the last two puck or last ten ball observations in the AirHockey and Table Tennis domains, respectively.
+We follow the standard training procedure from DDPM [35] to train all diffusion baselines.
+Model hyperparameters used for all domains are shown in Table I.
+We benchmark against the following LfD baselines: Behavioral cloning with Gaussian Mixture Models (BC-GMM) [52], Probabilistic Movement Primitives (ProMP) [26], and Action Chunking with Transformers (ACT) [19].
+We also compare against four diffusion baselines: Diffusion (Base), Motion Planning Diffusion (MPD1) [24], BEbhavior generation with ScOre-based Diffusion Policies (BESO) [41] and Movement Primitive Diffusion (MPD2) [42].
 Method Block Rate $\uparrow$ Time (s) T ms/step
 Imitation Learning Baselines
-Behavioral Cloning (BC) [49] 15.5% - - -
-ACT [19] 79.4% - - -
+Behavioral Cloning (BC-GMM) [52] 15.5% 0.001 - -
+ACT [19] 79.4% 0.011 - -
 Unconditioned Diffusion Models
-Unconstrained w/o Batch Filter 12.9% 0.159 20 7.95
-Unconstrained 62.6% 0.161 20 8.05
+Diffusion (Base) 12.9% 0.159 20 7.95
+Diffusion (Base) w/ Batch Filter 62.6% 0.161 20 8.05
 MPD1 [24] 63.5% 0.190 16 11.88
@@
 Puck Conditioned Diffusion Models
-Unconstrained w/o Batch Filter 46.2% 0.162 20 8.10
-Unconstrained 46.6% 0.163 20 8.15
+Diffusion (Base) 46.2% 0.162 20 8.10
+Diffusion (Base) w/ Batch Filter 46.6% 0.163 20 8.15
 MPD1 [24] 58.2% 0.191 16 11.94
@@
 KCGG (Ours) 47.3% 0.188 10 18.80
-TABLE II: Air-Hockey Defend Results: We compare our constraint sampling method against imitation learning and diffusion baselines. “T” is the number of diffusion timesteps.
+TABLE III: AirHockey Defend Results: We compare our constraint sampling method against imitation learning and diffusion baselines. “T” is the number of diffusion timesteps.
 VI-C Simulation Results: AirHockey
@@
 For each task, we test on 1000 initial start configurations.
-Our results for Hit (Table I) show that the diffusion model captures diverse motions best, achieving highest success rates for each shot type.
-Results for the Defend task (Table II), show KCGG enables unconditioned diffusion models to outperform all baselines, increasing hit rate by 18.9% over the best baseline.
+Our results for Hit (Table II) show that Diffusion (Base) captures diverse motions best, achieving highest success rates for each shot type.
+Results for the Defend task (Table III), show KCGG enables unconditioned diffusion models to outperform all baselines, increasing block rate by 18.9% over the best baseline.
 We use a fixed wall-clock time budget of 200 ms and tune sampling steps accordingly.
 Interestingly, KCGG does not improve puck-conditioned models.
-We hypothesize the conditioned base model overfits the score function to puck observations.
+We hypothesize that the conditioned model overfits the score function to puck observations.
+We conduct an experiment on Defend by using an obstacle avoidance constraint with KCGG instead, i.e.
+$c(\tau)=\min_{t\in[t_{s},t_{e}],q_{t}\in\tau}\frac{1}{\|F(q_{t})-b_{t}\|_{2}^{2}}$, and achieve a $2.9\%$ block rate, showing KCGG is compatible with other constraints.
 Figure 3: Comparison of Planned Trajectory: KCGG maintains a smooth joint angle profile in the constrained region (dotted lines) while constraint guidance introduces discontinuities.
-Figure 3 shows a comparison of the joint angle trajectories generated by KCGG and the original constraint guidance, revealing KCGG more effectively achieves the desired end-effector trajectory while adhering to the data distribution.
-The trajectory produced by the original constraint guidance exhibits significant inconsistencies in the constrained region whereas KCGG samples a smooth trajectory.
-VI-D Sampling Time Ablation
-We perform a sensitivity analysis comparing against other diffusion baselines by varying the number of diffusion timesteps (Figure 4).
-Despite requiring more time per step, KCGG outperforms diffusion baselines across all sampling time budgets, except for MPD2 with sampling time budget less than $70$ ms.
-Wall-clock time is evaluated as sampling step time vary for each method.
-All experiments were run on a single Nvidia GTX 1080.
-Figure 4: Sampling Speed vs Performance: We compare the Block Rate in Air-Hockey Defend across sampling times. KCGG scales to better performance with sampling time.
-As the number of diffusion timesteps increases, the performance of the baseline methods plateau.
-In contrast, KCGG scales effectively, reaching a block rate nearing 85%, while a method such as MPD2 does not, despite performing better with few sampling steps.
-Additional denoising steps allows KCGG to better adhere the trajectory samples to both the training data distribution and task constraints.
+Constraint wrt. $\tau_{i-1}$ Constraint wrt. $\tau_{0}$
+Gradient wrt. $\tau_{i-1}$ 63.5% (MPD1 [24]) 65.0%
+Gradient wrt. $\tau_{i}$ 70.0% 85.2% (Ours)
+TABLE IV: Success Rates reported ablating for design choices in KCGG in the AirHockey Defend domain
+Figure 3 shows a comparison of a joint angle trajectory generated by KCGG and MPD1 [24], revealing that the trajectory produced by MPD1 [24] exhibits significant inconsistencies in the constrained region whereas KCGG samples a smooth trajectory from the data distribution.
+Figure 4: Sampling Speed vs Performance: We compare the Block Rate in AirHockey Defend across sampling times. KCGG scales to better performance with sampling time.
 Left Push Left Lob Left Slice Left Topspin
@@
 ACT [19] 0% 0% 66% 64% 40% 88% 82% 14% 100% 70% 24% 88%
-Unconstrained Diffusion 28% 8% 72% 38% 6% 46% 10% 4% 12% 14% 10% 22%
+Diffusion (Base) 28% 8% 72% 38% 6% 46% 10% 4% 12% 14% 10% 22%
 MPD1 [24] 36% 22% 100% 70% 16% 100% 78% 8% 100% 52% 16% 100%
@@
 ACT [19] 56% 10% 90% 70% 26% 74% 28% 0% 34% 40% 22% 62%
-Unconstrained Diffusion 42% 6% 42% 36% 10% 40% 38% 0% 98% 50% 16% 82%
+Diffusion (Base) 42% 6% 42% 36% 10% 40% 38% 0% 98% 50% 16% 82%
 MPD1 [24] 66% 0% 94% 92% 18% 100% 84% 0% 100% 72% 0% 100%
@@
 KCGG (Ours) 100% 10% 100% 100% 32% 100% 72% 0% 100% 70% 22% 94%
-TABLE III: Table Tennis Results: We compare the hit rates (H), success rates (S), and correct stroke rates (CS) for four different shot types across two different ball launcher configurations (left and center launches).
+TABLE V: Table Tennis Results: We compare the hit rates (H), success rates (S), and correct stroke rates (CS) for four different shot types across two different ball launcher configurations (left and center launches) with 50 trials per condition.
+VI-D Ablation Studies
+Table IV reports ablation results for: (1) Calculating the gradient with respect to $\tau_{i-1}$ or $\tau_{0}$ and (2) Computing the constraint function with $\tau_{i-1}$ or $\tau_{0}$.
+We find that taking the gradient with respect to $\tau_{i}$ brings a small improvement of 6.5% over MPD1 and taking the constraint with respect to $\tau_{0}$ also provides a small 1.5% improvement.
+With both, our KCGG method achieves large gains over the baselines.
+We perform a sensitivity analysis against other diffusion baselines by varying the number of diffusion timesteps (Figure 4).
+Wall-clock time is evaluated as sampling step time vary for each method.
+All experiments were run on a single Nvidia GTX 1080.
+Despite requiring more time per step, KCGG outperforms diffusion baselines across all sampling time budgets, except for MPD2 with sampling time budget less than $70$ ms.
+As the number of diffusion timesteps increases, the performance of the baseline methods plateau.
+In contrast, KCGG scales, reaching a block rate near 85%, while a method such as MPD2 does not, despite performing better with few sampling steps.
+Additional denoising steps allows KCGG to better adhere the trajectory samples to both the training data distribution and task constraints.
 VI-E Real-world Results: Table Tennis
-We report results comparing diffusion models to imitation learning baselines in Table III with 50 samples per task.
-Expert demonstrations are used instead of traditional heuristics, which lack the flexibility to generate diverse table-tennis stroke profiles and are hard to design.
-Based on Air-Hockey Defend findings, we conditioned our models only on stroke type to avoid over-parameterizing the score function.
+Results comparing KCGG to all other baselines are reported in Table V (50 samples per task).
+Based on AirHockey findings, our models are conditioned only on the stroke type to avoid over-parameterizing the score function.
 In terms of task performance, KCGG shows an average 17.3% improvement in success rate across all tasks, outperforming or matching the highest baseline success rates in six of eight tasks.
 Specifically, KCGG surpasses prior diffusion constraint guidance methods (MPD1) with a notable 124.5% improvement.
-We also note a modest average 4% improvement in hit rate over the baselines, attributed mainly to high hit rates in the push and lob strokes.
+We also note a modest average 4% improvement in hit rate over the baselines.
 KCGG excels in most strokes but has a lower hit rate in slice and topspin.
@@
 KCGG consistently demonstrated the desired behavior in each task, shown by high correct stroke rates.
-Conversely, ACT, despite conditioning on stroke style, struggled to generate correct left/center strokes from ball observations, yielding a lower correct stroke rate.
+Conversely, ACT, despite conditioning on stroke style, struggled to generate the correct left/center strokes from ball observations, yielding a lower correct stroke rate.
 For BESO and MPD2, we observed that conditioning on ball observation history caused overfitting and collapse into certain stroke modalities (e.g.
-BESO mostly samples Center Lob strokes), illustrating the importance of constraint-guidance.
-We created stronger variants of these baselines by replacing ball history conditioning with ”ball oracle” conditioning, a categorical label indicating Left or Center launch distribution from the demonstration.
-This provides direct information on the ball launch location, which significantly improved BESO and MPD2 performance, removing mode collapse for a stronger comparison.
+BESO mostly samples Center Lob strokes), illustrating the importance of constraint guidance.
+We created stronger variants of these baselines by replacing ball history conditioning with “ball oracle” conditioning, a categorical label indicating Left or Center launch distribution from the demonstration.
+This provides direct information on the ball launch location, which significantly improved BESO and MPD2 performance.
 However, KCGG still outperformed these variants on average in hit and success rate.
@@
 VIII Conclusion
-Our work demonstrates KCGG’s effectiveness on small expert datasets in agile robotic tasks.
-Our algorithm is validated in both a simulated air hockey domain and on a real table tennis robot.
+Our work demonstrates KCGG’s effectiveness on small expert datasets in agile robotic tasks, through evaluations in both a simulated air hockey domain and on a real table tennis robot.
 Our real robot experiments show that we can learn diverse policies without access to a high-fidelity simulator.